OpenStudy (anonymous):

LIMIT FAST:

OpenStudy (anonymous):

Find the limit \[1+(-1)^n\]

OpenStudy (ipwnbunnies):

As n approaches what?

OpenStudy (sidsiddhartha):

limt n tends to what?

OpenStudy (anonymous):

guess what: infinity

OpenStudy (tkhunny):

Oh, a Smart Alexander. Nice. Try to be more respetcful and cooperative. What say you? Is "n" an integer? What does \((-1)^{n}\) do as n increases without bound?

OpenStudy (anonymous):

Find the limit of the row : \[x _{n}=1+(-1)^n\]

OpenStudy (tkhunny):

Much better. Using it as an index suggests that n is a Whole Number or Natural Number. Okay, now answer the question. What does \((−1)^{n}\) do as n increases without bound? How about as it takes on 1, 2, 3, and then 4? What does that piece do?

OpenStudy (anonymous):

I'm not that stupid @tkhunny \[(-1)^n=\left[\begin{matrix}-1 & if;n; isnot;even\\ 1 & if;n even\end{matrix}\right]\]

OpenStudy (anonymous):

that (-1) provokes me an undetermination in the limit: \[UNDETERMINATION:CASE:\left[ 1^{\infty } \right]\]

OpenStudy (tkhunny):

Who said you were stupid at all? There you have it. The whole expression flops around between 2 and 0. Therefore, the individual terms do not converge If we are talking about series, we can also say that there is no series convergence.